# Unveiling the Role of ChatGPT in Software Development: Insights from Developer-ChatGPT Interactions on GitHub

### v2

@@ Abstract.
The advent of Large Language Models (LLMs) has introduced a new paradigm in software engineering, with generative AI tools like ChatGPT gaining widespread adoption among developers. While ChatGPT’s potential has been extensively discussed, there is limited empirical evidence exploring its real-world usage by developers. This study bridges this gap by conducting a large-scale empirical analysis of ChatGPT-assisted development activities, leveraging a curated dataset, DevChat, comprising 2,547 unique shared ChatGPT links collected from GitHub between May 2023 and June 2024. Our study examines the characteristics of ChatGPT usage on GitHub (including the tendency, prompt turns distribution, and link descriptions) and identifies five categories of developers’ purposes for sharing developer-ChatGPT conversations during software development. Additionally, we analyzed the development-related activities where developers shared ChatGPT links to facilitate their workflows. We then established a mapping framework among data sources, activities, and SE tasks associated with these shared ChatGPT links. Our study offers a comprehensive view of ChatGPT’s application in real-world software development scenarios and provides a foundation for its future integration into software development workflows.
@@ 1. Introduction
The emergence of Large Language Models (LLMs) has fundamentally reshaped the landscape of Software Engineering (SE) practices (Hou et al., 2024). They introduce a novel paradigm for software development, offering capabilities that can potentially revolutionize the field. As one of the most typical and representative LLM-based tools, ChatGPT (OpenAI, ChatGPT, 2025), launched by OpenAI in November 2022, has sparked widespread interest within both academia and industry, leading to numerous studies exploring its applicability in software development, such as code generation and refinement (Sobania et al., 2023; Guo et al., 2024; Li et al., 2023), code summarization and comment generation (Sun et al., 2025; Su and McMillan, 2024), and program repair (Zhang et al., 2023).
@@
Notably, in May 2023, OpenAI released a new feature allowing users to share links within their ChatGPT interactions (ChatGPT Shared Links FAQ, 2025). This feature promotes the sharing of knowledge generated by ChatGPT and enhances team collaboration through shared developer-ChatGPT interactions. As a result, this feature has further boosted ChatGPT’s popularity among developers, surpassing other LLMs in developers’ communities. As evidenced by our trial search, before selecting ChatGPT as our target LLM, we conducted a preliminary search on GitHub for other LLMs that support shared conversation links, such as Gemini (Google, Gemini, 2025) , and found significantly fewer shared links. Consequently, we decided to use ChatGPT as the focus of this study, which is the most popular and widely used LLM on GitHub.
@@
Recently, Xiao et al. (Xiao et al., 2024) shared a dataset of developer-ChatGPT conversations used on GitHub, collected from July to October 2023, which ignited interest in studying ChatGPT’s practical use. However, due to its relatively small size and temporal coverage, the dataset is not sufficient to support our comprehensive understanding of the real-world prevalence and usability of ChatGPT in software development. While sharing ChatGPT links is a subset of ChatGPT’s overall usage in real-world practices, the links offer a valuable observational window into actual developers’ practices. Specifically, these shared instances of ChatGPT usage present cases that developers deemed the interaction noteworthy or useful enough to share, but to what extent the shared ChatGPT conversations impact the practical development is not the focus of this study. In this work, we mined and analyzed such shared and available ChatGPT links on GitHub to investigate the current landscape and patterns of ChatGPT usage in software development. For brevity, we henceforth refer to “publicly shared ChatGPT usage” simply as “the usage of ChatGPT” or “ChatGPT usage” in this paper.
@@
Additionally, compared to previous studies published in the Mining Challenge track of MSR 2024 (Mining Challenge Track of MSR 2024, 2024) that investigated the developer-ChatGPT interactions from certain perspectives (e.g., code generation, prompt interactions), our work comprehensively analyzed the usage of ChatGPT during the software development process on GitHub through those shared and available developer-ChatGPT interactions, including characterizing developer-ChatGPT interactions, uncovering developers’ motivations, analyzing the development-related activities developers use ChatGPT for, categorizing the concrete SE tasks performed with ChatGPT’s assistance. The integration of ChatGPT into software development workflows is still in its early stages, but it has already shown great promise in improving developer productivity and code quality (Vaillant et al., 2024). Therefore, our goal is to investigate how developers utilize ChatGPT in their daily routines to better anticipate its broader impact on the SE field.
@@
To address these gaps and achieve our goal, we conducted a large-scale empirical investigation into shared ChatGPT links, offering insights into LLM-assisted software development. Specifically, we first curated DevChat, a large-scale dataset that encompasses ChatGPT usage from May 2023 to June 17, 2024. DevChat contains 2,547 unique shared ChatGPT links related to a variety of development activities across five main development-related data sources on GitHub, i.e., Code, Issues, Commits, Pull Requests, and Discussions. Leveraging DevChat, we conducted an in-depth analysis of how developers utilize ChatGPT during software development.
@@
Our analysis reveals several key findings: Developers predominantly share ChatGPT links in Code (43.4%) and Commits (32.3%), and their enthusiasm for sharing ChatGPT links reached a peak in August 2023. Developers primarily employ ChatGPT for short, task-specific prompts, with two- or three-turn prompts being the most frequent. Contextual descriptions are prevalent in most shared ChatGPT links, particularly in Commits (99%), but are often absent in Code. In contrast, in Issues and Pull Requests, the absence of descriptions suggests an assumption of inherent content clarity. The developers’ purposes for sharing ChatGPT links on GitHub were identified and categorized into five categories, with Task Delegation being the dominant purpose, primarily to automate repetitive work. Moreover, we categorized the development-related activities within the shared ChatGPT conversations into two primary groups: Lifecycle Activities and Supporting Activities, with Software Implementation, Software Maintenance and Evolution emerged as the dominant activities. Furthermore, we established the mappings among the data sources, activities, and tasks of shared developer-ChatGPT conversations. These mappings provide a comprehensive overview of 39 specific SE tasks, with Code Generation & Completion and Code Modification & Optimization emerging as the most prevalent. These findings provide crucial empirical evidence for guiding future research and the development of AI-assisted tools, contributing to a deeper understanding of the evolving AI-empowered development paradigm.
@@
  • We curated DevChat, a large-scale dataset of developer-ChatGPT interactions relevant to software development on GitHub. Following deduplication and manual review, the cleaned dataset is available in our dataset (Li et al., 2025). This dataset enables future research into the practical applications of ChatGPT in software development.
  • We analyzed the characteristics of ChatGPT usage on GitHub (including the tendency, prompt turns distribution, and link descriptions). Additionally, we identified and categorized developers’ purposes for sharing developer-ChatGPT conversations during development in practice.
  • We examined the development-related activities for which developers shared ChatGPT links to support their workflows. Furthermore, we established a mapping relationship among data sources, activities, and SE tasks associated with these shared links, providing a comprehensive view of ChatGPT’s role in practical development scenarios.
@@ 2.1. Large Language Models for Software Engineering
LLMs are extensively utilized throughout various phases of the Software Development Lifecycle (SDLC), including requirements engineering, software design, software development, quality assurance, maintenance, and management (Hou et al., 2024). The most predominant phase is software development, and code generation and program repair being the most prevalent tasks for LLMs (Hou et al., 2024). For example, as a code generation tool empowered by LLM, GitHub Copilot (GitHub, Copilot, 2025) received much attention from software developers after its release. Vaithilingam et al. (Vaithilingam et al., 2022) conducted a user study with 24 participants on the usability of GitHub Copilot. They found that most participants preferred using Copilot, as Copilot provided a useful starting point for programming tasks and saved them the effort of searching code snippets online. Liang et al. (Liang et al., 2024) performed an exploratory study to investigate the usability of LLM-assisted programming tools, such as GitHub Copilot, with 410 programmers. Their results indicate that developers are motivated to use AI programming assistants (e.g., GitHub Copilot) because these tools help reduce keystrokes, speed up tasks, and recall syntax, but there is still a gap between developers’ needs and the tools’ output. Jin et al. (Jin et al., 2023) introduced an end-to-end program repair framework, InferFix, to fix critical security and performance bugs in Java and C#. By combining a static analyzer with a fine-tuned LLM, InferFix shows better performance than LLM baselines on patch generation in Java and C#. Moreover, a recent study of Poldrack et al. (Poldrack et al., 2023) reported the experimental results of AI-assisted coding with GPT4, including performing required coding tasks, refactoring, and improving the quality of existing code. Their results show that although AI coding tools are powerful, human involvement is still necessary to ensure the validity and accuracy of the results.
@@ 2.3. Conclusive Summary
Table 1 summarizes the key differences between our work and prior studies. Regarding the use of LLMs, prior studies (e.g., (Jin et al., 2023), (Poldrack et al., 2023), (Vaithilingam et al., 2022)) explored the characteristics of LLM-assisted programming, and it remains unclear which development-related activities pertain to leveraging LLMs and the characteristics of their usage. Regarding the use of ChatGPT, previous studies (e.g., (Aguiar et al., 2024), (Hao et al., 2024), (Sobania et al., 2023)) investigated the use of ChatGPT in software development practices, but they focused on specific topics such as ChatGPT-generated code, bug-fixing performance, and human-ChatGPT conversations, instead of providing a comprehensive analysis of ChatGPT usage during the software development lifecycle. Table 1 presents a comparison between our work and prior studies, highlighting the comprehensive perspective of our work.
Table 1. Comparison between our work and prior studies
Previous Studies Comparison with Our Work
Regarding the use of LLMs in SE (Hou et al., 2024), previous studies mainly focused on generation tasks (e.g., LLM-assisted programming (Liang et al., 2024)) and classification tasks (e.g., requirements classification (Luo et al., 2022)). Unlike previous empirical studies that primarily investigated classification and generation tasks (especially code generation), we attempted to uncover comprehensive development-related activities on leveraging LLMs and the characteristics of their usages across the software development lifecycle.
@@
Regarding the use of ChatGPT in software development, previous studies on the Mining Challenge Track of MSR 2024 (Mining Challenge Track of MSR 2024, 2024) primarily focused on the following aspects: Code Generation Quality and Developers’ Perceptions, such as analysis of the quality and correctness of ChatGPT-generated code (e.g., (Siddiq et al., 2024)), comparison between ChatGPT-generated code and human-written alternatives (e.g., (Rabbi et al., 2024)). Developers’ Behaviors and Practices, such as developers’ interactions with ChatGPT and practices in real-world scenarios (e.g., (Mohamed et al., 2024)). Code Integration and Modification: Integration and Modification of ChatGPT-generated code into projects (e.g., (AlOmar et al., 2024)). Prompt Interaction and Iterative Resolution: Investigation of the iterative nature of prompting and its use in issue resolutions (e.g., (Wu et al., 2024)). Compared to previous studies (AlOmar et al., 2024; Rabbi et al., 2024; Mohamed et al., 2024; Wu et al., 2024; Siddiq et al., 2024) that investigated the use of ChatGPT in specific aspects of software development, our work comprehensively analyzed the sharing of ChatGPT links during the software development process on GitHub. Specifically, we investigated practical ChatGPT sharing in real-world applications based on our established large and clean dataset (i.e., DevChat (Li et al., 2025)): Characterize how developers interact with ChatGPT (RQ1). Uncover the motivations behind the developer-ChatGPT interactions (RQ2). Analyze the development-related activities for which developers share ChatGPT links (RQ3). Categorize the concrete software engineering tasks performed by developers with the help of shared ChatGPT links (RQ4).
@@ 3.1. Research Questions
Rationale: RQ1 aims to characterize the nature and patterns of developers’ interactions with ChatGPT in real-world software development, such as the frequency and distribution of prompt turns within developer–ChatGPT interactions, and the contextual descriptions accompanying shared ChatGPT links. By analyzing these dimensions, RQ1 can uncover how developers leverage ChatGPT throughout the development process.
RQ2: Why do developers share ChatGPT links during software development?
Rationale: By exploring the diverse purposes for which developers utilize ChatGPT, RQ2 seeks to provide nuanced insights into the practical utility and limitations of ChatGPT in development, thereby laying the groundwork for more effective integration of Generative AI (GenAI) tools in software development workflows.
@@
Rationale: RQ3 is designed to map out a comprehensive landscape of ChatGPT usage during software development. By identifying and categorizing various development-related activities that are widely recognized in SWEBOK (Washizaki et al., 2024) and prior studies (Hou et al., 2024; Fan et al., 2023), this RQ quantitatively examines the distribution of ChatGPT links across these activities, offering insights into ChatGPT’s practical applications and impact on the software development process.

### v4

@@ Abstract.
The advent of Large Language Models (LLMs) has introduced a new paradigm in Software Engineering (SE), with generative AI tools like ChatGPT gaining widespread adoption among developers. While ChatGPT’s potential has been extensively discussed, empirical evidence about how developers actually use LLMs’ assistance in real-world practices remains limited. To bridge this gap, we conducted a large-scale empirical analysis of ChatGPT usage on GitHub, and we presented DevChat, a curated dataset of 2,547 publicly shared ChatGPT conversation links collected from GitHub between May 2023 and June 2024. Through comprehensively analyzing DevChat, we explored the characteristics of developer–ChatGPT interaction patterns and identified five key categories of developers’ purposes of using developer-ChatGPT conversations during software development. Additionally, we investigated the dominant development-related activities in which ChatGPT is used, and presented a mapping framework that links GitHub data sources, development-related activities, and SE tasks. The findings show that interactions are typically short and task-focused (most are 1-3 turns); developers share conversations mainly to delegate tasks, resolve problems, and acquire knowledge, revealing five purpose categories; ChatGPT is most frequently engaged for Software Implementation and Maintenance & Evolution; we identified 39 fine-grained SE tasks supported by ChatGPT, with Code Generation & Completion as well as Code modification & Optimization being the most prominent. Our study offers a comprehensive mapping of ChatGPT’s applications in real-world software development scenarios and provides a foundation for understanding LLMs’ practical roles in software development.
@@ 1. Introduction
The emergence of Large Language Models (LLMs) has fundamentally reshaped the landscape of Software Engineering (SE) practices (Hou et al., 2024). They introduce a novel paradigm for software development, offering capabilities that can potentially revolutionize the field. As one of the most typical and representative LLM-based tools, ChatGPT (OpenAI, ChatGPT, 2025), launched by OpenAI in November 2022, has sparked widespread interest within both academia and industry, leading to numerous studies exploring its applicability in software development, such as code generation and refinement (Sobania et al., 2023; Guo et al., 2024; Li et al., 2023), code summarization and comment generation (Sun et al., 2025; Su and McMillan, 2024), and program repair (Zhang et al., 2023). Although prior research has demonstrated ChatGPT’s capabilities and evaluated its performance on specific tasks, there remains limited empirical understanding of how developers actually use ChatGPT in real-world software development workflows, and for what purposes and tasks it is leveraged in daily practice.
@@
A key challenge in studying real-world ChatGPT usage is the lack of observable traces that reflect developers’ authentic practices. In May 2023, OpenAI released a new feature allowing users to share links within their ChatGPT interactions (ChatGPT Shared Links FAQ, 2025). This feature promotes the sharing of knowledge generated by ChatGPT and enhances team collaboration through shared developer-ChatGPT interactions. Developers create a unique record of ChatGPT-assisted development by embedding these links in GitHub workflows. As a result, this feature has further boosted ChatGPT’s popularity among developers, surpassing other LLMs in developers’ communities. As evidenced by our trial search, before selecting ChatGPT as our target LLM, we conducted a preliminary search on GitHub for other LLMs that support shared conversation links, such as Gemini (Google, Gemini, 2025) , and found significantly fewer shared links. Consequently, we decided to use ChatGPT as the focus of this study, which is the most popular and widely used LLM on GitHub.
@@
Recent work has begun to explore such developer–ChatGPT interactions. For example, Xiao et al. (Xiao et al., 2024) shared a dataset of developer-ChatGPT conversations used on GitHub, collected from July to October 2023, which ignited interest in studying ChatGPT’s practical use. However, due to its relatively small size and temporal coverage, the dataset is not sufficient to support our comprehensive understanding of the prevalence and usability of ChatGPT in software development. Compared to the developer survey of ChatGPT usage conducted by Vaillant et al. (Vaillant et al., 2024), our work provides empirical evidence of actual usage through observable ChatGPT links that developers proactively share. These links constitute concrete artifacts of ChatGPT usage, reflecting instances where developers considered the interactions sufficiently valuable to reference or reuse within development workflows. As such, our work offers complementary insights to perception- and recall-based survey studies by grounding the analysis in actually occurring usage data. We acknowledge that shared ChatGPT links represent only a subset of ChatGPT’s overall usage in practice. Nevertheless, the links offer a valuable observation window into real-world developers’ practices, capturing interactions that developers deem noteworthy or relevant enough to externalize. While this study does not aim to assess the direct impact of these shared conversations on development outcomes, it focuses on characterizing how ChatGPT is leveraged in developers’ workflows through these observable usage traces. In this work, we mined and analyzed such shared and available ChatGPT links on GitHub to investigate the current landscape and patterns of ChatGPT usage in software development. For brevity, we henceforth refer to “publicly shared ChatGPT usage” simply as “the usage of ChatGPT” or “ChatGPT usage” in this paper.
@@
Additionally, compared to previous studies published in the Mining Challenge track of MSR 2024 (Mining Challenge Track of MSR 2024, 2024) that investigated the developer-ChatGPT interactions from certain perspectives (e.g., code generation, prompt interactions), our work comprehensively analyzed the ChatGPT usage during the software development process on GitHub through the shared and available ChatGPT links, including characterizing developer-ChatGPT interactions, uncovering developers’ motivations, analyzing the development-related activities developers use ChatGPT for, mapping GitHub data sources, development-related activities, and concrete SE tasks performed with ChatGPT’s assistance. The integration of ChatGPT into software development workflows is still in its early stages, but it has already shown great promise in improving developer productivity and code quality (Vaillant et al., 2024).
@@
Therefore, our goal is to investigate how developers utilize ChatGPT in their daily routines to better anticipate its broader impact on the SE field. To this end, we conducted a large-scale empirical investigation into shared ChatGPT links, offering insights into LLM-assisted software development. Specifically, we presented DevChat, a curated dataset of 2,547 unique shared ChatGPT conversations collected from GitHub between May 2023 and June 2024. These conversations are drawn from five major GitHub data sources (i.e., Code, Issues, Commits, Pull Requests, and Discussions), capturing a diverse range of development scenarios. Leveraging DevChat, we conducted an in-depth analysis of how developers utilize ChatGPT during software development.
@@
Our analysis reveals several key findings: Developer–ChatGPT interactions are typically involving short, low-turn conversations (most are 1-3 turns); Developers’ purposes of using ChatGPT links on GitHub were identified and categorized into five categories, with Task Delegation being the dominant purpose; We categorized the development-related activities within the shared ChatGPT conversations into two primary groups: Lifecycle Activities and Supporting Activities, with Software Implementation, Software Maintenance and Evolution emerged as the dominant activities; By systematically mapping data sources, development-related activities, and fine-grained SE tasks from shared ChatGPT links, our analysis reveals that ChatGPT’s practical impact is predominantly concentrated in later SDLC phases, while its adoption in design tasks remains limited, with Code Generation & Completion and Code Modification & Optimization being the most prevalent. These findings offer valuable empirical evidence for guiding future research and the development of AI-assisted tools, contributing to a deeper understanding of the evolving AI-empowered development paradigm.
@@
  • Dataset for Investigating Shared ChatGPT Links Usages: We curated DevChat, a dataset comprising 2,547 unique shared ChatGPT conversations on GitHub. Following deduplication and manual review, the cleaned dataset is available at [27]. This dataset enables future research into the practical applications of ChatGPT in software development.
  • Empirical Evidence of ChatGPT Usages on GitHub: We presented the characteristics of ChatGPT usages on GitHub (RQ1); We identified and categorized developers’ purposes of using developer-ChatGPT conversations during development in practice (RQ2); We examined the development-related activities for which developers shared ChatGPT links to support their development workflows (RQ3); We established a mapping relationship among GitHub data sources, development-related activities, and SE tasks associated with these shared ChatGPT links, providing a comprehensive view of ChatGPT’s role in practical development scenarios (RQ4).
  • Insights for Practitioners and Researchers: Based on our findings, we distilled concrete, evidence-based insights for practice and research, including context-aware integration of ChatGPT into development workflows, prioritizing tool support for dominant usage scenarios, and identifying underexplored SE areas where ChatGPT adoption remains limited. These insights provide evidence-based guidance for the responsible and effective use of ChatGPT.
@@ 2.1. Large Language Models for Software Engineering
LLMs are extensively utilized throughout various phases of the Software Development Lifecycle (SDLC), including requirements engineering, software design, software development, quality assurance, maintenance, and management (Hou et al., 2024). The most predominant phase is software development, and code generation and program repair being the most prevalent tasks for LLMs (Hou et al., 2024). For example, as a code generation tool empowered by LLM, GitHub Copilot (GitHub, Copilot, 2025) received much attention from software developers after its release. Vaithilingam et al. (Vaithilingam et al., 2022) conducted a user study with 24 participants on the usability of GitHub Copilot. They found that most participants preferred using Copilot, as Copilot provided a useful starting point for programming tasks and saved them the effort of searching code snippets online. Sergeyuk et al. (Sergeyuk et al., 2025) presents an empirical study on the practical use of AI-based coding assistants, examining developers’ adoption patterns, perceptions, and challenges. Their findings highlight usage practices and limitations of AI-based coding assistants, offering insights that complement research on the broader impacts of LLM-based tools. Liang et al. (Liang et al., 2024) performed an exploratory study to investigate the usability of LLM-assisted programming tools, such as GitHub Copilot, with 410 programmers. Their results indicate that developers are motivated to use AI programming assistants (e.g., GitHub Copilot) because these tools help reduce keystrokes, speed up tasks, and recall syntax, but there is still a gap between developers’ needs and the tools’ output. Jin et al. (Jin et al., 2023) introduced an end-to-end program repair framework, InferFix, to fix critical security and performance bugs in Java and C#. By combining a static analyzer with a fine-tuned LLM, InferFix shows better performance than LLM baselines on patch generation in Java and C#. Moreover, a recent study of Poldrack et al. (Poldrack et al., 2023) reported the experimental results of AI-assisted coding with GPT4, including performing required coding tasks, refactoring, and improving the quality of existing code. Their results show that although AI coding tools are powerful, human involvement is still necessary to ensure the validity and accuracy of the results.
@@ 2.3. Conclusive Summary
Table 1 provides a structured overview of prior studies, the existing research gaps, and the distinctions between our work and previous efforts. The table is organized by four columns: Theme (topics of related work), Previous Studies (summarizations of previous studies), Gap (what remains unexplored), and Comparison (how our study fills these gaps).
While previous studies (see Table 1) have investigated the capabilities of LLMs and ChatGPT in software development, their focuses have largely been on task-specific evaluations, isolated interaction patterns, or model-centric performance metrics. Consequently, there is limited understanding of how developers leverage ChatGPT into everyday software development workflows and the broader range of activities supported by ChatGPT in practice.
Our study addresses these gaps by adopting a developer-centric perspective and analyzing real-world usage of ChatGPT. Our work provides a more comprehensive and nuanced understanding of ChatGPT usage across the software development lifecycle, revealing practical usage patterns, underlying developers’ motivations, and specific types of SE tasks it supports.
Table 1. Gap and Comparison between our work and prior studies
Theme Previous Studies Gap Comparison
The use of LLMs in SE LLMs have been used for various SE tasks, including automated code generation (Jiang et al., 2026; Sergeyuk et al., 2025), code completion (Husein et al., 2025), bug detection and repair (Zhang et al., 2023), and documentation synthesis (Dvivedi et al., 2024), demonstrating LLMs’ potential to augment developer productivity and in software development. Previous studies primarily adopt task-centric perspectives, focusing on model capabilities, performance comparisons, or qualitative assessments of output quality. While they provide valuable insights into what LLMs can do for specific SE tasks, comprehensive insights into how developers actually leverage LLMs into their daily development workflows remain unexplored. Our work complements and extends existing LLM4SE research by shifting the focus from model-centric capabilities to developer-centric practices, providing a holistic empirical foundation for understanding the practical ChatGPT usage, SE task distribution assisted by ChatGPT, and potential areas for future AI-assisted development tool enhancements.
@@
The use of ChatGPT in software development As one of the most representative LLMs, ChatGPT has been used in diverse SE tasks. For example, previous studies on the Mining Challenge Track of MSR 2024 (Mining Challenge Track of MSR 2024, 2024) primarily focused on four aspects: (1) Code Generation Quality and Developers’ Perceptions, such as analysis of the quality and correctness of ChatGPT-generated code (e.g., (Siddiq et al., 2024)), comparison between ChatGPT-generated code and human-written alternatives (e.g., (Rabbi et al., 2024)). (2) Developers’ Behaviors and Practices, such as developers’ interactions with ChatGPT and practices in real-world scenarios (e.g., (Mohamed et al., 2024)). (3) Code Integration and Modification: Integration and Modification of ChatGPT-generated code into projects (e.g., (AlOmar et al., 2024)). (4) Prompt Interaction and Iterative Resolution: Investigation of the iterative nature of prompting and its use in issue resolutions (e.g., (Wu et al., 2024)). While previous studies provide valuable insights into specific use of ChatGPT, they remain limited in scope, focusing on isolated tasks or interaction patterns rather than covering the comprehensive lifecycle of software development. In particular, prior work does not systematically analyze how developer-ChatGPT interactions reflect developers’ practical usage across diverse development activities, their underlying motivations behind the developer-ChatGPT interactions, or the specific SE tasks facilitated by ChatGPT. Our work comprehensively analyzed the sharing of ChatGPT links during the software development process on GitHub. Specifically, we investigated practical ChatGPT sharing in real-world applications based on our established large and cleaned dataset (i.e., DevChat (Li et al., 2025)): Characterize how developers interact with ChatGPT (RQ1). Uncover the motivations behind the developer-ChatGPT interactions (RQ2). Analyze the development-related activities for which developers share ChatGPT links (RQ3). Categorize the concrete software engineering tasks performed by developers with the help of shared ChatGPT links (RQ4).
@@ 3.1. Research Questions
Rationale: RQ1 aims to analyze the distributional and usage characteristics of shared ChatGPT links on GitHub, such as the frequency and distribution of prompt turns within developer–ChatGPT interactions, and the contextual descriptions accompanying shared ChatGPT links. By analyzing these dimensions, RQ1 can uncover how developers leverage ChatGPT throughout the development process.
RQ2: What are the purposes for which developers use ChatGPT based on the shared ChatGPT links?
Rationale: By examining the specific purposes for which developers use ChatGPT based on the shared ChatGPT links during software development, RQ2 seeks to provide nuanced insights into the developers’ intentions in employing ChatGPT beyond the technical tasks involved. This purpose-oriented understanding complements task-level analysis and provides insight into how ChatGPT shapes collaboration and information exchange in software development.
@@
Rationale: RQ3 is designed to map out a comprehensive landscape of ChatGPT usage during software development. By identifying and categorizing various development-related activities that are widely recognized in SWEBOK (Washizaki and others, 2024) and prior studies (Hou et al., 2024; Fan et al., 2023), this RQ quantitatively examines the distribution of ChatGPT links across these activities, offering insights into ChatGPT’s practical applications and impact on the software development process.